Find the equation of the line through (2, -3) and perpendicular to the line defined by the equation 4x + 5y + 6 = 0
Answer by Cromlix(4381)

4x + 5y + 6 = 0
Sort into y = mx + c form.
5y = -4x - 6
y = -4/5x - 6/5
Lines that are perpendicular to one
another have slopes (m) that multiply
together to give -1
m1 * m2 = -1
y = -4/5x - 6/5
m1 = -4/5
m2 will equal 5/4
-4/5 * 5/4 = -1
Using equation y - b = m(x - a)
a = 2, b = -3, m = 5/4
y -(-3) = 5/4(x - 2)
y + 3 = 5/4x - 10/4
y = 5/4x - 10/4 + 12/4 (3 = 12/4)
y = 5/4x + 2/4
OR multiply thro' by 4
4y = -5x + 2
Hope this helps. :-)